Ten years from now, A will be twice as old as B. Five years ago, A was three times as old as B. What are the present ages of A and B?
Answer by scott8148(6628)

"Ten years from now, A will be twice as old as B" ___ A + 10 = 2(B + 10)

"Five years ago, A was three times as old as B" ___ A - 5 = 3(B - 5)

subtract the equations (to eliminate A); solve for B; then substitute back to find A